Let K be a field and f : Z → K the homomorphism of

integers into K.

a) Show that the kernel of f is a prime ideal. If f is an embedding,

then we say that K has characteristic zero.

b) If kerf f= {0}, show that kerf is generated by a prime number

p. In this case we say that K has characteristic p.


1
Expert's answer
2022-06-23T18:01:49-0400

Recall that kernel of any ring homomorphism is an ideal. Hence "Ker(f)" is an ideal of "\mathbb{Z}". Then the ideal "Ker(f)" is a prime ideal if and only if "\mathbb{Z}/Ker(f)" is an integral domain. Indeed if "xKer(f),yKer(f)" are in "\mathbb{Z}/Ker(f)" and "xKer(f)yKer(f)=Ker(f)". Then "xy\in Ker(f)", then "f(xy)=f(x)f(y)". But "K" is a field either "f(x)=0" or "f(y)=0". i.e. either "x\in Ker(f)" of "y\in Ker(f)". Hence "\mathbb{Z}/Ker(f)" is an integral domain.


On the other hand if "Ker(f)=\{0\}", then "f(n)\ne 0" for any "n\ne0". In particular "0\ne f(1)=f(1)f(1)" and "f(1)" is an element of the field implies multiplying from right by its inverse we have "f(1)" is the multiplicative identity of the field "K". Hence for any integer n and any non-zero element "a\in K" we have "na=f(n)\ne 0" as product of two non-zero element in a field is non-zero. Hence characteristic of the field "K" is zero.


(b) On the other hand since "\mathbb{Z}/Ker(f)" is an integral domain when "Ker(f)\ne0", the ideal "Ker(f)" must be a prime ideal. But the only prime ideals of the ring "\mathbb{Z}" are the ones generated by prime numbers. Hence "\mathbb{Z}/Ker(f)\cong\mathbb{Z}_p". Hence "K" has a smallest subfield isomorphic to "\mathbb{Z}_p". It follows that "K" has characteristic "p" for the prime "p".